# Resisting Backdoor Attacks in Federated Learning via Bidirectional Elections and Individual Perspective

### v1

@@ 5.2 Experimental Setup
Approaches are implemented with PyTorch 1.10 (Paszke et al. 2019). The code will be online-available after blinded review. For all approaches, we build a network with 2 convolutional layers followed by 2 fully-connected (FC) layers on MNIST, Fashion MNIST and FEMNIST, a network with 6 convolutional layers followed by 1 FC layer on CIFAR-10, and a GRU layer followed by 1 FC layer on Sent140. Detailed model backbones are available in Appendix B.1. For Snowball, we build a simple VAE with three layers, and set $M$ as $\frac{K}{2}$, $\check{M}=0.1K$, $M^{E}=0.05K$ on FEMNIST and otherwise $0.04K$, $E^{VI}$ and $E^{VT}$ higher than 270 and 30, respectively. Detailed hyperparameters are listed in Appendix C. These models are trained by stochastic gradient descendant (SGD) optimizer with learning rate started at 0.01 and decays by 0.99 after each round.
@@
Approaches are evaluated by backdoor task accuracy (BA) and main task accuracy (MA). MA is the best accuracy of global model on the test set among all rounds since in reality there may be a validation set. BA is the probability that the global model identifies the test samples with triggers as the target class of the attack in the round where the highest MA is achieved.
@@ 5.3 Performance
The performance of Snowball and its peers is presented in Table 1 and 2, where the best BA among realistic approaches is marked in bold. Each value is averaged on three runs with different random seeds. For FLIP, we leave the results in CIFAR-10, FEMNIST and Sent140 blank since we have tried but always encountered with NaN problem even in the official implementation with the global model replaced by ours.
It is shown that Snowball is effective in defending against backdoor attacks on all five datasets, showing a competitive BA with Ideal, while existing approaches either fails to effectively withstand backdoor attacks or significantly degrades MA. Due to the large number of attackers and unclear boundary between benign and infected model updates, Krum and RLR struggle to distinguish between them. Although CRFL and FLAME have stronger ability to resist backdoor attacks than Krum and RLR, their MA decreases due to the DP noise. FLDetector fails to can defend against backdoor attacks because it fails to trace the history model updates of a client due to partial participation. DnC is effective on MNIST but fails in other complex scenarios, since it is based on a subset of model parameters. If the intersection between the subset for detection and the small number of parameters affected by backdoor attacks is not large, DnC may be ineffective. FLIP is effective on MNIST and Fashion MNIST, but it causes severe decrease on MA the same as in (Zhang et al. 2023).
@@
As in (Nguyen et al. 2022), we test Snowball with different PDR to show the resilience of Snowball against attacks of varying strengths. When PDR is high, one wrongly included infected update can cause catastrophic consequences. We select some of the baselines and two representative datasets, i.e., MNIST and FEMNIST with label skew and feature skew, respectively. As in Figures 6(a) and 6(b), Snowball can effectively defend against backdoor attacks on data with different PDR. We have also noticed that FLAME gradually fails to defend attacks as PDR increases, since the noise may be not enough to disturb stronger attacks. DnC performs well with high PDR since more model parameters will be affected there, increasing the likelihood of affected parameters being sampled by the down-sampling.
Limited by space, more evaluations are in Appendix D.
@@ 5.4 Hyperparameter Sensitivity of VAE
Hyperparameters of VAE in top-down election are easy be appropriately set since Snowball is not sensitive to them. Figure 7(a) presents BA of Snowball with different combinations of $E^{VI}$ and $E^{VT}$. Generally, larger $E^{VI}$ and $E^{VT}$ would not make BA worse, since the VAE can be trained better. As shown in Figure 7(a), if they are too small, VAE underfits and fails to distinguish between benign and infected model updates. $S^{H}$ is the dimensionality of hidden layer outputs of the encoder and decoder of the VAE, and $S^{L}$ is that of the latent feature $\mathbf{z}$ generated by the encoder, respectively. As shown in Figure 7(b), Snowball does not exhibit significant differences in BA within their selected range of values, indicating that they are easy to set to appropriate values.
@@ 6 Conclusion
This work focuses on defending against backdoor attacks in FL and proposes a novel approach named Snowball. It enables an individual perspective that treats each model update as an agent electing model updates for aggregation, and conducts bidirectional election to select models to be aggregated, i.e., a) bottom-up election where each model update votes to several peers such that a few model updates are elected as selectees for aggregation; and b) top-down election, where selectees progressively enlarge themselves focusing on differences between model updates. Extensive experiments conducted on five real-world datasets demonstrate the superior performance of Snowball to resist backdoor attacks compared with SOTA approaches when dealing with the situations in which 1) the non-IIDness of data is complex and the PDR is not high such that the benign and infected model updates do not obviously gather in different positions, and 2) the ratio of attackers to all clients is not low. Besides, Snowball is easy to be integrated into existing FL systems.

### v2

@@ 5.2 Experimental Setup
Approaches are implemented with PyTorch 1.10 (Paszke et al. 2019). For all approaches, we build a network with 2 convolutional layers followed by 2 fully-connected (FC) layers on MNIST, Fashion MNIST and FEMNIST, a network with 6 convolutional layers followed by 1 FC layer on CIFAR-10, and a GRU layer followed by 1 FC layer on Sent140. Detailed model backbones are available in Appendix B.1. For Snowball, we build a simple VAE with three layers, and set $M$ as $\frac{K}{2}$, $\check{M}=0.1K$, $M^{E}=0.05K$ on FEMNIST and otherwise $0.04K$, $E^{VI}$ and $E^{VT}$ higher than 270 and 30, respectively. Detailed hyperparameters are listed in Appendix C. These models are trained by the stochastic gradient descendant (SGD) optimizer with a learning rate starting at 0.01 and decays by 0.99 after each round.
@@
Approaches are evaluated by backdoor task accuracy (BA) and main task accuracy (MA). MA is the best accuracy of the global model on the test set among all rounds since in reality there may be a validation set. BA is the probability that the global model identifies the test samples with triggers as the target class of the attack in the round where the highest MA is achieved.
@@ 5.3 Performance
The performance of Snowball and its peers is presented in Table 1 and 2, where the best BA among realistic approaches is marked in bold. Each value is averaged on three runs with different random seeds. For FLIP, we leave the results in CIFAR-10, FEMNIST and Sent140 blank since we have tried but always encountered NaN problem even in the official implementation with the global model replaced by ours.
It is shown that Snowball is effective in defending against backdoor attacks on all five datasets, showing a competitive BA with Ideal, while existing approaches either fail to effectively withstand backdoor attacks or significantly degrade MA. Due to the large number of attackers and unclear boundaries between benign and infected model updates, Krum and RLR struggle to distinguish between them. Although CRFL and FLAME have a stronger ability to resist backdoor attacks than Krum and RLR, their MA decreases due to the DP noise. FLDetector fails to defend against backdoor attacks because it fails to trace the history model updates of a client due to partial participation. DnC is effective on MNIST but fails in other complex scenarios, since it is based on a subset of model parameters. If the intersection between the subset for detection and the small number of parameters affected by backdoor attacks is not large, DnC may be ineffective. FLIP is effective on MNIST and Fashion MNIST, but it causes a severe decrease in MA the same as in (Zhang et al. 2023).
@@
Following Nguyen et al. (2022), we test Snowball with different PDR to show its resilience to attacks of varying strengths. When PDR is high, one wrongly included infected update can cause catastrophic consequences. We select some of the baselines and two representative datasets with label skew and feature skew, respectively. As in Figures 6, Snowball can effectively defend against backdoor attacks on data with different PDR. We have also noticed that FLAME gradually fails to defend against attacks as PDR increases, since the noise may be not enough to disturb stronger attacks. DnC performs well with high PDR since more model parameters will be affected there, increasing the likelihood of affected parameters being sampled by the down-sampling.
Limited by space, more experimental evaluations are left in Appendix D.
@@ 5.4 Hyperparameter Sensitivity of VAE
Hyperparameters of VAE in Snowball are easy be set since the VAE is not sensitive to them. Figure 7(a) presents BA of Snowball with different combinations of $E^{VI}$ and $E^{VT}$. Generally, larger $E^{VI}$ and $E^{VT}$ would not make BA worse, since the VAE can be trained better. But if they are too small, the VAE underfits and fails to distinguish between benign and infected model updates. $S^{H}$ is the dimensionality of hidden layer outputs of the encoder and decoder of the VAE, and $S^{L}$ is that of the latent feature $\mathbf{z}$ generated by the encoder, respectively. As shown in Figure 7(b), Snowball does not exhibit significant differences in BA in the selected range of values, showing that they are easy to set to appropriate values.
@@ 6 Conclusion
This work proposes a novel approach named Snowball for defending against backdoor attacks in FL. It enables an individual perspective that treats each model update as an agent electing model updates for aggregation, and conducts bidirectional election to select models to be aggregated, i.e., a) bottom-up election where each model update votes to several peers such that a few model updates are elected as selectees for aggregation; and b) top-down election, where selectees progressively enlarge themselves focusing on differences between model updates. Experiments conducted on five real-world datasets demonstrate the superior resistance to backdoor attacks of Snowball compared to SOTA approaches in situations where 1) the non-IIDness of data is complex and the PDR is not high such that the benign and infected model updates do not obviously gather in different positions, and 2) the ratio of attackers to all clients is not low. Besides, Snowball can be easily integrated into existing FL systems.
Acknowledgments
This work was supported in part by the Key Research Project of Zhejiang Province under Grant 2022C01145 and in part by the National Science Foundation of China under Grants 62125206 and U20A20173.